### v1

# A Conceptual Framework For White Box Neural Networks

Abstract
This paper introduces semantic features as a general conceptual framework for fully explainable neural network layers. A well-motivated proof of concept model for relevant subproblem of MNIST consists of 4 such layers with the total of ~4.8K learnable parameters. The model is easily interpretable, achieves human-level adversarial test accuracy with no form of adversarial training, requires little hyperparameter tuning and can be quickly trained on a single CPU. The general nature of the technique bears promise for a paradigm shift towards radically democratised and truly generalizable white box neural networks. The code is available at https://github.com/314-Foundation/white-box-nn.
1 Introduction
Current state of the art (SOTA) neural networks are large, unwieldy black boxes offering little insight into their decision-making process, require prohibitive amounts of resources and are prone to adversarial attacks not seen during training [10, 3]. The latter implies that models rely on unintended features shared between train and test, giving false sense of generalisability [4, 6]. These issues drastically limit the usability of current generation of neural networks, especially in critical domains such as healthcare, cybersecurity or autonomous systems. Yet animal brains apparently learn rich and robust patterns in mostly unsupervised sensory input - even brains as simple as drosophila brain with ~135k neurons. This discrepancy indicates a need for radically simplified AI.
This paper aims to lay foundation for a new paradigm for training models by introducing a conceptual framework for building white box neural networks. The emphasis is put on theoretical and qualitative aspects of proposed solutions and less on tweaking quantitative metrics. Following this spirit it’s important to train a proof of concept (PoC) network on Minimum Viable Dataset (MVD) - problem easy enough to get rid of all the obfuscating details but still hard enough to be relevant. MNIST dataset is falsely considered to be "solved" or "too easy" while SOTA models are hardly adequate in terms of their complexity, lack interpretability and are vulnerable to adversarial perturbations not seen during training [9]. This is true not only for the entire 10-digit dataset but even for certain pairs of digits [11] and therefore the MVD for this paper is chosen to be the binary subset of MNIST consisting of images of "3" and "5" (one of the most often confused pairs), which actually might be even harder than the entire MNIST in terms of average adversarial accuracy. This is not to say that similar results cannot be obtained on more complex datasets - preliminary experiments suggest that it is exactly the opposite - but in fact one of the goals of such selection of MVD is to make a clear and concise argument for the effectiveness and potential of presented method.
The structure of the paper is as follows:
  • Section 2. gives a rigorous definition of semantic feature and argues for the generality of the notion;
  • Section 3. builds a carefully motivated PoC 4-layer white box neural network for selected MVD;
  • Section 4. analyses the trained model both qualitatively and quantitatively in terms of explainability, reliability and adversarial robustness.
The paper ends with ideas for further research and acknowledgments.
2 Semantic features
In machine learning inputs are represented as tensors. Yet the standard Euclidean topology in tensor space usually fails to account for small domain-specific variations of features. For example shifting an image by 1 pixel is semantically negligible but usually results in a distant output in $L_{2}$ metric.
This observation inspires the following definition:
Definition:
A semantic feature $f_{\mathbf{P}\mathbf{L}}$ of dimension $s\in\mathbb{N}$ is a tuple $(f,\mathbf{P},\mathbf{L})$ where:
  • $f$ base, $f\in\mathbb{R}^{s}$,
  • $\mathbf{P}$ parameter set, $\mathbf{P}\subset\mathbb{R}^{r}$ for some $r\in\mathbb{N}$
  • $\mathbf{L}$ locality function : $\mathbf{P}\rightarrow\textrm{Aut}(\mathbb{R}^{s},\mathbb{R}^{s})$ where Aut is the family of differentiable automorphisms of $\mathbb{R}^{s}$
Additionally the feature $(f,\mathbf{P},\mathbf{L})$ is called differentiable iff $\mathbf{L}$ is differentiable (on some neighbourhood of $\mathbf{P}$).
Intuitively a semantic feature consists of a base vector $f$ and a set of it’s "small" variations $f_{p}=\mathbf{L}(p)(f)$ for all $p\in\mathbf{P}$; these variations may be called poses after [8]. Thus $(f,\mathbf{P},\mathbf{L})$ is locality-sensitive in the adequate topology capturing semantics of the domain. The differentiability of $\mathbf{L}(p)$ and of $\mathbf{L}$ itself allow for gradient updates of $f$ and $\mathbf{P}$ respectively. The intention is to learn both $f$ and $\mathbf{P}$ while defining $\mathbf{L}$ explicitly as an inductive bias111by the ”no free lunch” theorem a general theory of learning systems has to adequately account for the inductive bias. for the given modality.
2.1 Examples of semantic features
To better understand the definition consider the following examples of semantic features:
  1. 1.
    real-valued $f_{\mathbf{P}\mathbf{L}}$
    • $f\in\mathbb{R}$
    • $\mathbf{P}$ a real interval $[p_{min},p_{max}]$ for $p_{min}<=0<=p_{max}$
    • $\mathbf{L}$ function that maps a real number $p$ to the function $g\mapsto g+p$ for $g\in\mathbb{R}$
  2. 2.
    convolutional $f_{\mathbf{P}\mathbf{L}}$
    • $f$ square 2D convolutional kernel of size $(c_{out},c_{in},k,k)$
    • $\mathbf{P}$ set of $2\times 2$ rotation matrices
    • $\mathbf{L}$ function that maps a rotation matrix to the respective rotation of a 2D kernel222can be made differentiable as in [7, 5].
  3. 3.
    affine $f_{\mathbf{P}\mathbf{L}}$
    • $f$ vector representing a small 2D spatial feature (e.g. image patch containing a short line)
    • $\mathbf{P}$ set of $2\times 3$ invertible augmented matrices333preferably close to the identity matrix.
    • $\mathbf{L}$ function that maps a $2\times 3$ matrix to the respective 2D affine transformation of images2
  4. 4.
    xor $f_{\mathbf{P}\mathbf{L}}$
    • $f$ single-entry vector of dimension $n\in\mathbb{N}$
    • $\mathbf{P}\subset\{0,1,\ldots,n-1\}$
    • $\mathbf{L}$ function that maps a natural number $p$ to the function that rolls the non-zero element of $g$ by $p$ coordinates to the right, where $g$ is a single-entry vector of dimension $n$
  5. 5.
    logical $f_{\mathbf{P}\mathbf{L}}$
    • $f$ concatenation of $k$ single-entry vectors and some dense vectors in-between
    • $\mathbf{P}\subset\{0,1,\ldots,n_{0}-1\}\times\ldots\times\{0,1,\ldots,n_{k-1}-1\}$
    • $\mathbf{L}$ function that maps $p\in\mathbf{P}$ to the respective rolls of single-entry vectors (leaving the dense vectors intact)
2.2 Matching semantic features
The role of semantic features is to be matched against the input to uncover the general structure of the data. Suppose that we have already defined a function match: $(\mathbb{R}^{c},\mathbb{R}^{s})\rightarrow\mathbb{R}$ that measures an extent to which datapoint $d\in\mathbf{R}^{c}$ contains a feature $g\in\mathbb{R}^{s}$. Now let’s define
$\mathrm{SFmatch}(d,(f,\mathbf{P},\mathbf{L}))=\max_{p\in\mathbf{P}}\mathrm{match}(d,\mathbf{L}(p)(f))$ (1)
If $\mathbf{L}$ captures domain topology adequately then we may think of matching $f_{\mathbf{P}\mathbf{L}}$ as a form of local inhibition along $\mathbf{L}_{\mathbf{P}}(f)=\{L(p)(f),p\in P\}$. Therefore semantic feature defines a XOR gate over $\mathbf{L}_{\mathbf{P}}(f)$. On the other hand every $f_{p}\in\mathbf{L}_{\mathbf{P}}(f)$ can be viewed as a conjunction (AND gate) of its non-zero coordinates444i.e. $f_{p}$ is a conjunction of lower-level features.. Therefore semantic features turn out to be a natural way of expressing logical claims about the classical world - in the language that can be "understood" by neural networks. This is especially apparent in the context of logical semantic features which allow to explicitly express logical sentences, e.g. sentence $(A\vee B)\wedge\neg C\wedge\neg D$ can be expressed by a logical $f_{\mathbf{P}\mathbf{L}}$ where:
  • $f=[1,0,-1,-1]$
  • $\mathbf{P}=\{0,1\}$
  • $\mathbf{L}(0)(f)=[1,0,-1,-1]$ and $\mathbf{L}(1)(f)=[0,1,-1,-1]$
Semantic features capture the core characteristic of any semantic entity - having many possible states but being at exactly one state at a time. As the remainder of the paper will show this turns out to be a sufficiently strong regularization - backpropagation through appropriately chosen $\mathbf{L}$ results in easily interpretable $f$ and $\mathbf{P}$.
It remains to characterise the match function. Usually it can be defined as the scalar product $d\cdot emb(g)$ where $emb:\mathbb{R}^{s}\rightarrow\mathbb{R}^{c}$ is some natural embedding of the feature space to the input space. In particular for our previous examples lets set the following:
  • for real-valued $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=\mathrm{int}(d==g)$ where $d,g\in\mathbb{R}$
  • for convolutional $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=\max_{0<=i<c_{out}}(d*\frac{g^{i}}{||g^{i}||_{2}})$ where $g$ is a 2D convolutional kernel of shape $(c_{out},c_{in},k,k)$, $g^{i}$ is the $i$-th convolutional filter of $g$, $d$ is an image patch of shape $(c_{in},k,k)$ 555alternatively we could treat every $g^{i}$ as a separate semantic feature but will we stick to the max option for simplicity.
  • for affine $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=d\cdot\frac{g}{||g||_{2}}$ where $g$ is a 2D image of shape $(c_{in},k,k)$ and $d$ is a 2D image patch of the same shape
  • for logical $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=d\cdot g$ for $g,d\in\mathbb{R}^{c}$
The occasional $L_{2}$ normalization in the above examples is to avoid favouring norm-expanding $\mathbf{L}(p)$.
3 Network architecture
In this section we will use semantic features to build a PoC white box neural network model that classifies MNIST images of "3" and "5" in a transparent and adversarially robust way. We will build the network layer by layer arguing carefully for every design choice. Note that the network construction is not an abstract process and must reflect the core characteristics of the chosen dataset. There is no free lunch, if we want to capture the semantics of the dataset we need to encode it in the architecture - more or less explicitly. The framework of semantic features allows to do this in a pretty natural way. Note that a single neural network layer will consist of many parallel semantic features of the same kind.
The network consists of the following 4 layers stacked sequentially. Note that for simplicity we don’t add bias to any of those layers. Te resulting model has around $4.8$K parameters.
3.1 Two Step Layer
MNIST datapoint is a $28\times 28$ grayscale image. Its basic building block is a pixel $x\in[0,1]$. Despite being allowed to assume any value between 00 and $1$ it is actually (semantically) a ternary object - it’s either ON, OFF or MEH666the MEH state is an undecided state in-between.. Therefore the semantic space should squash the $[0,1]$ interval into those 3 values.
A real-valued $f_{\mathbf{P}\mathbf{L}}$ with match and $\mathbf{L}$ defined as in the previous section identifies numbers in certain interval around $f$. This means that a layer consisting of multiple real-valued $f_{\mathbf{P}\mathbf{L}}$, provided that the relevant intervals don’t overlap, can be represented as a single locally-constant real-valued function. This allows us to simplify the implementation of such layer and instead of learning the constrained parameters of multiple $f_{\mathbf{P}\mathbf{L}}$ we can learn a single parameterised real-valued function that serves as an entire layer. In order to group pixel intensities into 3 abstract values we can draw inspiration from the $\mathrm{softsign}(x)=\frac{x}{1+x}$ function or rather from it’s parameterised version $\mathrm{softsign}(x,t,s)=\frac{x-t}{s+x}$ and “glue” two of such step functions together to obtain (Parameterised) Two Step function shown in Figure 2. The exact implementation of the Two Step function is provided in Appendix A. We initialize the layer setting init_scales_div=10.
3.2 Convolutional Layer
A pixel-wise function is not enough to classify pixel as ON or OFF - the semantics of our MVD require a pixel to exist in a sufficiently bright region to be considered as ON. Therefore the next layer will consist of a single convolutional semantic feature $(f,\mathbf{P},\mathbf{L})$:
  • $f$ 2D kernel of shape $(2,1,5,5)$ (concatenation of $(g^{0},g^{1})$, both of shape $(1,1,5,5)$)
  • $\mathbf{P}$ fixed set of $k$ distinct rotations by a multiple of the $\frac{360^{\circ}}{k}$ angle, for $k=32$
  • $\mathbf{L}$ as defined earlier
The layer performs $\mathrm{SFmatch}(d(x,y),f_{\mathbf{P}\mathbf{L}})$ with every pixel $(x,y)$ of image $d$. This means that it matches the two rotated filters $g^{0}$ and $g^{1}$ with $(x,y)$ and takes the maximum across all the matches777a single number, not a pair of numbers, as in the chosen definition of match for convolutional $f_{\mathbf{P}\mathbf{L}}$.. Intuitively this layer checks if the pixel has a sufficiently bright neighbourhood.
The layer is followed by ReLU activation. We initialize $g^{0}$ and $g^{1}$ as the identity kernel of shape $(1,1,5,5)$ plus Gaussian noise $\sim\mathcal{N}(0,0.1)$.
3.3 Affine Layer
The basic semantic building blocks of MNIST digits are fragments of various lines together with their rough locations in the image (same line at the top and the bottom often has different semantics). In short they are shapes at locations. The semantic identity of shape is not affected by small888close to the identity augmented matrix affine transformations. The next shape-extracting layer will therefore consist of 8 affine semantic features $(f,\mathbf{P},\mathbf{L})$ defined as follows:
  • $f$ 2D image patch of shape $(1,20,20)$
  • $\mathbf{P}$ set of 32 affine $2\times 3$ augmented matrices
  • $\mathbf{L}$ as defined earlier
For simplicity every $f_{\mathbf{P}\mathbf{L}}$ in this layer is located in the center of $28\times 28$ image, i.e. we zero-pad $f$ on every side before applying the affine transformations. Intuitively this matches the image against a localized shape in a way that is robust to small affine perturbations of the shape.
The layer is followed by ReLU activation. We initialize affine matrices as $2\times 2$ identity matrix plus Gaussian noise $\sim\mathcal{N}(0,0.01)$ concatenated with $1\times 2$ matrix filled with Gaussian noise $\sim\mathcal{N}(0,1)$.
3.4 Logical Layer
After Affine Layer has extracted predictive shapes it remains to encode logical claims such as "number 3 is A or B and not C and not D" etc. Since the previous layer is learnable we can enforce a preferable structure on the input neurons to the Logical Layer. This layer will consist of 2 logical $f_{\mathbf{P}\mathbf{L}}$ (one for every label) defined as follows:
  • $f$ vector of dimension 8
  • $\mathbf{P}=\{0,1,2,3\}$
  • $\mathbf{L}$ as defined earlier
We initialize $f^{0}=[\frac{4}{10},0,0,0,-\frac{1}{10},-\frac{1}{10},-\frac{1}{10},-\frac{1}{10}]$ and $f^{1}=[-\frac{1}{10},-\frac{1}{10},-\frac{1}{10},-\frac{1}{10},\frac{4}{10},0,0,0,]$. The exact initialization values are less important then their sign and relative magnitude, however they should not be too large as we will train the network using CrossEntropyLoss.
4 Results
4.1 Training
The model was trained to optimize the CrossEntropyLoss for 15 epochs with Adam optimizer, learning rate $3e{-3}$, weight decay $3e{-6}$. The only augmentation was the Gaussian noise $\sim\mathcal{N}(0.05,0.25)$ applied with $0.7$ probability999mean is slightly positive because of the sparsity of data - subtracting pixel value is on average more harmful than adding it.. Those parameters were chosen "by hand" after few rounds of experimentation. The results seem pretty robust to the selection of hyperparameters as long as they are within reasonable bounds. Training a single epoch on a single CPU takes around 9s.
Reliability curveRefer to caption
Learning curveRefer to caption
Figure 1: (left) Reliability test-time curve for the classic 40-step PGD Attack. (right) Learning curve.
4.2 Quantitive results
We don’t bother measuring the clean test metrics as it is a flawed approach that tells little about the true generalisation capability of the model. Every metric reported here is computed under strong adversarial regime.
The model achieves ~92% accuracy under AutoAttack[2] (with default parameters: norm=’inf’, eps=0.3, step_size=0.1). The metrics are almost identical for the much faster 40-step PGD Attack. Samples are usually misclassified with low confidence and therefore we compute a reliability curve - a precision-recall curve where the positive class is defined as the set of correctly classified samples (this means that the precision for 100% recall is exactly equal to the accuracy). It turns out that for 80% recall model achieves human-level 98% adversarial precision (Figure 1). This means that in practice human intervention would be required only for the easily detected 20% of adversarial samples to achieve human level reliability of the entire system (and most real-life samples are not adversarial).
4.3 Layer inspection
Lets inspect the trained model layers.
4.3.1 Two Step Layer
Figure A show the initial and learned shapes of the Two Step function. Te function smoothly thresholds the input at $0.365$ and then at $0.556$ which means that the interval $[0,0.365]$ is treated as OFF, $[0.365,0.556]$ as MEH and $[0.556,1]$ as ON. Bear in mind that this is not the final decision regarding the pixel state - it’s just the input to the next layer.
Figure 2: Two Step layer - initial and learned.
4.3.2 Convolutional Layer
Quick inspection of the Figure 3 shows that the Convolutional Layer has indeed learned a meaningful filters that check if the central pixel lays inside an adequately structured region - intuitively the matching value corresponds to the degree the model considers the pixel as being ON.
Figure 3: Learned convolutional kernels in 32 predefined poses.
4.3.3 Affine Layer
Figure 4 shows the learned affine base features. Figure B in the Appendix B shows learned features in 32 learned feature-specific poses. It’s hard to imagine substantially more meaningful features.
Figure 4: Learned affine base features.
4.3.4 Logical Layer
The two learned logical base features are as follows:
  • $f^{0}=[1.0700,0.0000,0.0000,0.0000,-0.7800,-0.7100,-0.8000,-0.8100]$
  • $f^{1}=[-0.8100,-0.7600,-0.8600,-0.8000,1.0500,0.0000,0.0000,0.0000]$
This is exactly as expected - $f_{\mathbf{P}\mathbf{L}}^{0}$ defines number "3" as an entity having large SFmatch with any of the first 4 affine semantic features and small SFmatch with all of the remaining 4 affine semantic features. For $f_{\mathbf{P}\mathbf{L}}^{1}$ its the other way round.
4.4 Decision boundary
Figure 5 shows the minimum perturbations required to change model’s predictions under Boundary Attack[1] for sample test images. The added perturbations are easily spotted and understood by humans. Note that by definition model is extremely unconfident on those samples and in practice the perturbations required to fool the entire system would have to be much larger.
Figure 5: Boundary Attack: minimal perturbations that flip model prediction (model confidence at around 50%).
Figure 6: Visualization of test data as "seen" by the first 3 consecutive layers.
Figure 6 shows how the data is "seen" by the first 3 layers. The consecutive rows and lables denote respectively:
  1. 1.
    adversarial input and predicted class probabilities
  2. 2.
    output of Two Step Layer
  3. 3.
    output of Convolutional Layer
  4. 4.
    top pose of affine features predicting label "3" (with the match value) [^10] These rows are actually overlaid with the middle row to make it easier to see what has been matched.
  5. 5.
    top pose of affine features predicting label "5" (with the match value)10
Note that in Figure 6 there are 3 misclassified examples and for every one of them it’s easy to understand why the model made the mistake.
5 Significance and further research
The importance of explainability is hard to overestimate. An automatic system discovering meaningful predictive patterns in data could significantly offload the task of finding the underlying causal structure of observed phenomena. This has a great potential to impact science, diagnostics, big data and many other fields. Combined with the demonstrated level of efficiency the resulting democratisation could open up exciting areas for growth worldwide, favouring diverse open source solutions and providing considerable level of robustness against centralisation.
The general nature of semantic features makes it reasonable to expect that extending these results to more complex datasets and domains should be relatively straightforward - but may require some honest engineering. The obvious ideas for further work include:
  • use semantic features in self-supervised setting to obtain interpretable dimensionality reduction;
  • study more complex logical semantic features in more elaborate scenarios;
  • figure out a way to make logical semantic features differentiable, i.e. make $\mathbf{P}$ learnable;
  • design richer modes of spatial variation, e.g. hierarchical affine semantic features to capture variations of compound objects (objects consisting of affine parts are themselves affine if you "zoom out" appropriately);
  • find adequate semantic topology for color space;
  • integrate affine features with sliding window approach, i.e. implement weight sharing of similar features at distant locations;
  • define semantic features for sound (inspiration for semantic invariants could be drawn from music theory);
  • define semantic features for text, e.g. to capture semantic invariance of permutations of words;
6 Acknowledgments
This research has been entirely self-funded - I saved money, left my job and took the risk of following independently a depressingly underfunded and underestimated line of research.
However it needs to be noted that I had the opportunity to do preliminary research on adversarial robustness as a half-time activity during my employment at MIM Solutions. Special thanks go to Piotr Sankowski, who introduced me to the issue of adversarial vulnerability, and both to Piotr Sankowski and Piotr Wygocki who - as my former employers - decided to trust me with that research as a half-time part of my contract. It was precisely thanks to this opportunity that I developed a strong intuition that modern neural network architectures are inherently flawed and need to be fundamentally re-examined.
The story of this paper - should the results be recognized by the community - indicates a need for deep systemic and cultural changes that would empower individuals and smaller institutions to allow them to afford putting values and long-term informed risks over short-term profits. This is especially needed in larger public institutions like universities. Otherwise we might get stuck following "easy" directions which leads to a lethal combination of demoralisation and deterioration of quality.
Appendix A Implementation of Two Step Layer
class TwoStepFunction(Module):
    """Simplifiedreal-valuedSFLayerforinputsin[0,1]"""
    def init_weights(self):
        self.a0 = nn.Parameter(torch.tensor(0.5))
        self.a1 = nn.Parameter(torch.tensor(0.5))
        self.t0 = nn.Parameter(torch.tensor(0.25))
        self.t1 = nn.Parameter(torch.tensor(0.75))
        self.scales = nn.Parameter(
            (torch.ones((4,)) / self.init_scales_div)
    def softsign(self, x, threshold, scale, internal):
        a = self.a1 if threshold > 0 else self.a0
        x = x - threshold
        x = x / (scale + x.abs())
        if internal:
            shift = threshold
            shift = shift / (scale + shift.abs())
            x = x + shift
            x = a * x / shift.abs()
            return x
        div = x.abs().max() / (1 - a)
        x = x / div
        x = x + threshold.sgn() * a
        return x
    def transform(self, x):
        return 2 * x - 1
    def forward(self, x):
        x = self.transform(x)
        t0, t1 = self.transform(self.t0), self.transform(self.t1)
        params = [
            (t0, self.scales[0], False),
            (t0, self.scales[1], True),
            (t1, self.scales[2], True),
            (t1, self.scales[3], False),
        xs = [self.softsign(x, *p) for p in params]
        masks = [
            (x < t0).float(),
            (t0 <= x).float() * (x < 0.0).float(),
            (0.0 <= x).float() * (x < t1).float(),
            (t1 <= x).float(),
        x = sum(m * xx for (m, xx) in zip(masks, xs))
        return x
Appendix B Learned Affine Features With Learned Poses

### v2

@@
-A Conceptual Framework For White Box Neural Networks
+Towards White Box Deep Learning
 Abstract
-This paper introduces semantic features as a general conceptual framework for fully explainable neural network layers.
-A well-motivated proof of concept model for relevant subproblem of MNIST consists of 4 such layers with the total of ~4.8K learnable parameters.
-The model is easily interpretable, achieves human-level adversarial test accuracy with no form of adversarial training, requires little hyperparameter tuning and can be quickly trained on a single CPU.
-The general nature of the technique bears promise for a paradigm shift towards radically democratised and truly generalizable white box neural networks.
+This paper introduces semantic features as a candidate conceptual framework for building inherently interpretable neural networks.
+A proof of concept model for informative subproblem of MNIST consists of 4 such layers with the total of ~5K learnable parameters.
+The model is well-motivated, inherently interpretable, requires little hyperparameter tuning and achieves human-level adversarial test accuracy - with no form of adversarial training!
+These results and the general nature of the approach warrant further research on semantic features.
 The code is available at https://github.com/314-Foundation/white-box-nn.
 1 Introduction
-Current state of the art (SOTA) neural networks are large, unwieldy black boxes offering little insight into their decision-making process, require prohibitive amounts of resources and are prone to adversarial attacks not seen during training [10, 3].
-The latter implies that models rely on unintended features shared between train and test, giving false sense of generalisability [4, 6].
-These issues drastically limit the usability of current generation of neural networks, especially in critical domains such as healthcare, cybersecurity or autonomous systems.
-Yet animal brains apparently learn rich and robust patterns in mostly unsupervised sensory input - even brains as simple as drosophila brain with ~135k neurons.
-This discrepancy indicates a need for radically simplified AI.
-This paper aims to lay foundation for a new paradigm for training models by introducing a conceptual framework for building white box neural networks.
-The emphasis is put on theoretical and qualitative aspects of proposed solutions and less on tweaking quantitative metrics.
-Following this spirit it’s important to train a proof of concept (PoC) network on Minimum Viable Dataset (MVD) - problem easy enough to get rid of all the obfuscating details but still hard enough to be relevant.
-MNIST dataset is falsely considered to be "solved" or "too easy" while SOTA models are hardly adequate in terms of their complexity, lack interpretability and are vulnerable to adversarial perturbations not seen during training [9].
-This is true not only for the entire 10-digit dataset but even for certain pairs of digits [11] and therefore the MVD for this paper is chosen to be the binary subset of MNIST consisting of images of "3" and "5" (one of the most often confused pairs), which actually might be even harder than the entire MNIST in terms of average adversarial accuracy.
-This is not to say that similar results cannot be obtained on more complex datasets - preliminary experiments suggest that it is exactly the opposite - but in fact one of the goals of such selection of MVD is to make a clear and concise argument for the effectiveness and potential of presented method.
-The structure of the paper is as follows:
+Machine learning systems used in safety-critical applications such as healthcare, military, autonomous vehicles or cybersecurity need to be reliable, i.e.
+they cannot fail unpredictably.
+The black box nature of state of the art (SOTA) deep neural networks (DNN) make it exceedingly hard to give such guarantees, which limits their usability to non-critical domains.
+The recent research in Explainable AI (XAI) focuses predominantly on post-hoc attribution methods which may provide explanations but, due to the complexity of SOTA, are hard to be converted to reliable fail-safes.
+This is a burning issue because of the wide susceptibility of DNNs to distribution shifts and adversarial attacks [18, 5] meaning that models can make arbitrarily incorrect predictions and their reported high test-time metrics don’t guarantee robust real-life performance111models often rely on spurious features that leak from test to train [6]..
+The recently popular Neurosymbolic AI [17] consists of higher-level methods that focus on augmenting neural network representations (machine perception) with background structured knowledge (reasoning) but don’t address the potential failures of the former.
+Existing methods of improving DNN robustness are far from ideal.
+Adversarial Training [11] being one of the strongest empirical defenses against attacks [1] is slow, data-hungry [15, 8] and results in robust generalisation gap [13, 21] as models overfit to the train-time adversaries.
+Thus, the biggest challenge of empirical approaches is reliable evaluation of robustness as more and more sophisticated attacks break the previously successful defenses [3, 4].
+To address these challenges the field of certified robustness has emerged with the goal of providing theoretically certified upper bounds for adversarial error [9].
+However the exact certification is NP-Complete [20] and therefore has serious limitations in practice as it’s limited to relatively small problems and small networks and/or to probabilistic approaches.
+Another shortcoming is that such guarantees can only be given for a precisely defined threat model which is usually assumed to be a $L_{p}$-bounded perturbation.
+The methods, results and theoretical obstacles can be vastly different222for example SOTA certified robustness for MNIST is ~94% for $L_{\infty},eps=0.3$ and only ~73% for $L_{2},eps=1.58$, see https://sokcertifiedrobustness.github.io/leaderboard/.
+for different $p\in\{0,1,2,\ldots,\infty\}$ and often require training the model specifically towards chosen $L_{p}$-certification.
+The highlighted issues are relevant for datasets as simple as MNIST [16] or even some of its binary subsets [19] - despite being "toy" and almost linearly separable problems, they still remain elusive in terms of robustness, rising further questions about the true generalisation capabilities of SOTA as no neural network model can be honestly considered reliable and adequate333in terms of internal complexity; SOTA models define inappropriately complicated functions.
+for this allegedly "simple" problem.
+1.1 Proposed approach
+This paper proposes an intermediate approach to DNN reliability and introduces a candidate theoretical framework for building white box neural networks.
+Building and analysing a white box DNN would be similar to writing and debugging a procedural program - it doesn’t have to come with theoretical certificates but needs to be easily understandable and internalizable by the programmer.
+A good software has to be designed well enough to account for reasonable failure scenarios, handle foreseeable exceptions and, should an unforseen bug happen, be amenable for fixes.
+If the design and inner mechanics of a DNN are sufficiently clear then working with such model wouldn’t be very different from maintaining a standard application.
+Current SOTA models are large, unwieldy black boxes that require prohibitive amounts of resources; yet animal brains apparently learn rich and robust patterns in mostly unsupervised sensory input - even brains as simple as drosophila brain with ~135K neurons!
+This scale seems easily maintainable with proper abstractions.
+After all, the core algorithms used today are mostly unchanged since ~30 years!
+What if there really exists a better way?
+To find it we must start with the basics and take the agile approach.
+Focusing on the simplest relevant problem cuts off the obfuscating details and speeds up the experimentation phase.
+With that in mind the Minimum Viable Dataset (MVD) for this paper is chosen to be the binary subset of MNIST consisting of images of "3" and "5" - as this is one of the most often confused pairs under adversaries [19] and therefore might be considered harder than the entire MNIST in terms of average adversarial accuracy444the preliminary results on the full MNIST are discussed in Appendix C..
+1.2 The core idea
+This paper introduces semantic features as a general idea for sharing weights inside a neural network layer.
+The rigorous definition will be given in the next section.
+The concept is similar to that of "inverse rendering" in Capsule Networks [14] where features have many possible states and the best-matching state has to be found.
+Identifying different states by the subsequent layers gives rise to controlled dimensionality reduction.
+Thus semantic features aim to capture the core characteristic of any semantic entity - having many possible states but being at exactly one state at a time.
+This is in fact a pretty strong regularization.
+As shown in this paper, choosing appropriate layers of semantic features for the MVD results in what can be considered as a white box neural network.
+The remainder of the paper is structured as follows:
   • Section 2. gives a rigorous definition of semantic feature and argues for the generality of the notion;
   • Section 3. builds a carefully motivated PoC 4-layer white box neural network for selected MVD;
-  • Section 4. analyses the trained model both qualitatively and quantitatively in terms of explainability, reliability and adversarial robustness.
+  • Section 4. analyses the trained model both qualitatively and quantitatively in terms of explainability, reliability, adversarial robustness and ablation;
+  • Section 5. provides a short discussion of the results;
 The paper ends with ideas for further research and acknowledgments.
@@
 The differentiability of $\mathbf{L}(p)$ and of $\mathbf{L}$ itself allow for gradient updates of $f$ and $\mathbf{P}$ respectively.
-The intention is to learn both $f$ and $\mathbf{P}$ while defining $\mathbf{L}$ explicitly as an inductive bias111by the ”no free lunch” theorem a general theory of learning systems has to adequately account for the inductive bias.
+The intention is to learn both $f$ and $\mathbf{P}$ while defining $\mathbf{L}$ explicitly as an inductive bias555by the ”no free lunch” theorem a general theory of learning systems has to adequately account for the inductive bias.
 for the given modality.
@@
     • $\mathbf{P}$ set of $2\times 2$ rotation matrices
-    • $\mathbf{L}$ function that maps a rotation matrix to the respective rotation of a 2D kernel222can be made differentiable as in [7, 5].
+    • $\mathbf{L}$ function that maps a rotation matrix to the respective rotation of a 2D kernel666can be made differentiable as in [12, 7].
   3. 3.
@@
     • $f$ vector representing a small 2D spatial feature (e.g. image patch containing a short line)
-    • $\mathbf{P}$ set of $2\times 3$ invertible augmented matrices333preferably close to the identity matrix.
-    • $\mathbf{L}$ function that maps a $2\times 3$ matrix to the respective 2D affine transformation of images2
+    • $\mathbf{P}$ set of $2\times 3$ invertible augmented matrices777preferably close to the identity matrix.
+    • $\mathbf{L}$ function that maps a $2\times 3$ matrix to the respective 2D affine transformation of images6
   4. 4.
@@
     • $\mathbf{P}\subset\{0,1,\ldots,n-1\}$
-    • $\mathbf{L}$ function that maps a natural number $p$ to the function that rolls the non-zero element of $g$ by $p$ coordinates to the right, where $g$ is a single-entry vector of dimension $n$
+    • $\mathbf{L}$ function that maps a natural number $p$ to the function that rolls elements of $g\in\mathbb{R}^{n}$ by $p$ coordinates to the right
   5. 5.
@@
 $\mathrm{SFmatch}(d,(f,\mathbf{P},\mathbf{L}))=\max_{p\in\mathbf{P}}\mathrm{match}(d,\mathbf{L}(p)(f))$ (1)
-If $\mathbf{L}$ captures domain topology adequately then we may think of matching $f_{\mathbf{P}\mathbf{L}}$ as a form of local inhibition along $\mathbf{L}_{\mathbf{P}}(f)=\{L(p)(f),p\in P\}$.
-Therefore semantic feature defines a XOR gate over $\mathbf{L}_{\mathbf{P}}(f)$.
-On the other hand every $f_{p}\in\mathbf{L}_{\mathbf{P}}(f)$ can be viewed as a conjunction (AND gate) of its non-zero coordinates444i.e.
+If $\mathbf{L}$ captures domain topology adequately then we may think of matching $f_{\mathbf{P}\mathbf{L}}$ as a form of local inhibition along $\mathbf{L}_{\mathbf{P}}(f)=\{L(p)(f),p\in P\}$ and in this sense semantic feature defines a XOR gate over $\mathbf{L}_{\mathbf{P}}(f)$.
+On the other hand every $f_{p}\in\mathbf{L}_{\mathbf{P}}(f)$ can be viewed as a conjunction (AND gate) of its non-zero coordinates888i.e.
 $f_{p}$ is a conjunction of lower-level features..
-Therefore semantic features turn out to be a natural way of expressing logical claims about the classical world - in the language that can be "understood" by neural networks.
-This is especially apparent in the context of logical semantic features which allow to explicitly express logical sentences, e.g.
+Thus semantic features turn out to be a natural way of expressing logical claims about the classical world - in the language that can be used by neural networks.
+This is particularly apparent in the context of logical semantic features which can express logical sentences explicitly, e.g.
 sentence $(A\vee B)\wedge\neg C\wedge\neg D$ can be expressed by a logical $f_{\mathbf{P}\mathbf{L}}$ where:
-  • $f=[1,0,-1,-1]$
-  • $\mathbf{P}=\{0,1\}$
-  • $\mathbf{L}(0)(f)=[1,0,-1,-1]$ and $\mathbf{L}(1)(f)=[0,1,-1,-1]$
-Semantic features capture the core characteristic of any semantic entity - having many possible states but being at exactly one state at a time.
-As the remainder of the paper will show this turns out to be a sufficiently strong regularization - backpropagation through appropriately chosen $\mathbf{L}$ results in easily interpretable $f$ and $\mathbf{P}$.
+  • $f=[1,0,-1,-1]$;           $\mathbf{P}=\{0,1\}$;           $\mathbf{L}(0)(f)=[1,0,-1,-1]$ and $\mathbf{L}(1)(f)=[0,1,-1,-1]$.
 It remains to characterise the match function.
@@
   • for real-valued $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=\mathrm{int}(d==g)$ where $d,g\in\mathbb{R}$
-  • for convolutional $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=\max_{0<=i<c_{out}}(d*\frac{g^{i}}{||g^{i}||_{2}})$ where $g$ is a 2D convolutional kernel of shape $(c_{out},c_{in},k,k)$, $g^{i}$ is the $i$-th convolutional filter of $g$, $d$ is an image patch of shape $(c_{in},k,k)$ 555alternatively we could treat every $g^{i}$ as a separate semantic feature but will we stick to the max option for simplicity.
+  • for convolutional $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=\max_{0<=i<c_{out}}(d*\frac{g^{i}}{||g^{i}||_{2}})$ where $g$ is a 2D convolutional kernel of shape $(c_{out},c_{in},k,k)$, $g^{i}$ is the $i$-th convolutional filter of $g$, $d$ is an image patch of shape $(c_{in},k,k)$ 999alternatively we could treat every $g^{i}$ as a separate semantic feature but will we stick to the max option for simplicity.
   • for affine $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=d\cdot\frac{g}{||g||_{2}}$ where $g$ is a 2D image of shape $(c_{in},k,k)$ and $d$ is a 2D image patch of the same shape
@@
 Note that a single neural network layer will consist of many parallel semantic features of the same kind.
-The network consists of the following 4 layers stacked sequentially. Note that for simplicity we don’t add bias to any of those layers. Te resulting model has around $4.8$K parameters.
+The network consists of the following 4 layers stacked sequentially. Note that for simplicity we don’t add bias to any of those layers. Te resulting model has around $5$K parameters.
 3.1 Two Step Layer
@@
 Its basic building block is a pixel $x\in[0,1]$.
-Despite being allowed to assume any value between 00 and $1$ it is actually (semantically) a ternary object - it’s either ON, OFF or MEH666the MEH state is an undecided state in-between..
+Despite being allowed to assume any value between 00 and $1$ it is actually (semantically) a ternary object - it’s either ON, OFF or MEH101010the MEH state is an undecided state in-between..
 Therefore the semantic space should squash the $[0,1]$ interval into those 3 values.
@@
 The layer performs $\mathrm{SFmatch}(d(x,y),f_{\mathbf{P}\mathbf{L}})$ with every pixel $(x,y)$ of image $d$.
-This means that it matches the two rotated filters $g^{0}$ and $g^{1}$ with $(x,y)$ and takes the maximum across all the matches777a single number, not a pair of numbers, as in the chosen definition of match for convolutional $f_{\mathbf{P}\mathbf{L}}$..
+This means that it matches the two rotated filters $g^{0}$ and $g^{1}$ with $(x,y)$ and takes the maximum across all the matches111111a single number, not a pair of numbers, as in the chosen definition of match for convolutional $f_{\mathbf{P}\mathbf{L}}$..
 Intuitively this layer checks if the pixel has a sufficiently bright neighbourhood.
@@
 In short they are shapes at locations.
-The semantic identity of shape is not affected by small888close to the identity augmented matrix affine transformations.
+The semantic identity of shape is not affected by small121212close to the identity augmented matrix affine transformations.
 The next shape-extracting layer will therefore consist of 8 affine semantic features $(f,\mathbf{P},\mathbf{L})$ defined as follows:
@@
   • $\mathbf{L}$ as defined earlier
-For simplicity every $f_{\mathbf{P}\mathbf{L}}$ in this layer is located in the center of $28\times 28$ image, i.e.
-we zero-pad $f$ on every side before applying the affine transformations.
+For simplicity every $f_{\mathbf{P}\mathbf{L}}$ in this layer is zero-padded evenly on every side and identified with the resulting $28\times 28$ image.
 Intuitively this matches the image against a localized shape in a way that is robust to small affine perturbations of the shape.
@@
 The model was trained to optimize the CrossEntropyLoss for 15 epochs with Adam optimizer, learning rate $3e{-3}$, weight decay $3e{-6}$.
-The only augmentation was the Gaussian noise $\sim\mathcal{N}(0.05,0.25)$ applied with $0.7$ probability999mean is slightly positive because of the sparsity of data - subtracting pixel value is on average more harmful than adding it..
+The only augmentation was Gaussian noise $\sim\mathcal{N}(0.05,0.25)$ applied with $0.7$ probability131313mean is slightly positive because of the sparsity of data - subtracting pixel value is on average more harmful than adding it..
 Those parameters were chosen "by hand" after few rounds of experimentation.
@@
 Training a single epoch on a single CPU takes around 9s.
-Reliability curveRefer to caption
-Learning curveRefer to caption
+Reliability curve
+Learning curve
 Figure 1: (left) Reliability test-time curve for the classic 40-step PGD Attack. (right) Learning curve.
-4.2 Quantitive results
+4.2 Quantitative results
 We don’t bother measuring the clean test metrics as it is a flawed approach that tells little about the true generalisation capability of the model.
@@
 4.3.1 Two Step Layer
-Figure A show the initial and learned shapes of the Two Step function.
-Te function smoothly thresholds the input at $0.365$ and then at $0.556$ which means that the interval $[0,0.365]$ is treated as OFF, $[0.365,0.556]$ as MEH and $[0.556,1]$ as ON.
+Figure A shows the initial and learned shapes of the Two Step function.
+The function smoothly thresholds input at $0.365$ and then at $0.556$ which means that the interval $[0,0.365]$ is treated as OFF, $[0.365,0.556]$ as MEH and $[0.556,1]$ as ON.
 Bear in mind that this is not the final decision regarding the pixel state - it’s just the input to the next layer.
@@
 4.3.2 Convolutional Layer
-Quick inspection of the Figure 3 shows that the Convolutional Layer has indeed learned a meaningful filters that check if the central pixel lays inside an adequately structured region - intuitively the matching value corresponds to the degree the model considers the pixel as being ON.
+Quick inspection of the Figure 3 shows that the Convolutional Layer has indeed learned meaningful filters that check if the central pixel lays inside an adequately structured region - intuitively the matching value corresponds to the degree the model considers the pixel as being ON.
 Figure 3: Learned convolutional kernels in 32 predefined poses.
@@
 Figure 6: Visualization of test data as "seen" by the first 3 consecutive layers.
-Figure 6 shows how the data is "seen" by the first 3 layers. The consecutive rows and lables denote respectively:
+Figure 6 shows how the data is "seen" by the first 3 layers. The consecutive rows and labels denote respectively:
   1. 1.
@@
   4. 4.
-    top pose of affine features predicting label "3" (with the match value) [^10] These rows are actually overlaid with the middle row to make it easier to see what has been matched.
+    top pose of affine features predicting label "3" (with the match value)141414these rows are actually overlaid with the middle row to make it easier to see what has been matched.
   5. 5.
-    top pose of affine features predicting label "5" (with the match value)10
+    top pose of affine features predicting label "5" (with the match value)14
 Note that in Figure 6 there are 3 misclassified examples and for every one of them it’s easy to understand why the model made the mistake.
-5 Significance and further research
-The importance of explainability is hard to overestimate.
-An automatic system discovering meaningful predictive patterns in data could significantly offload the task of finding the underlying causal structure of observed phenomena.
-This has a great potential to impact science, diagnostics, big data and many other fields.
-Combined with the demonstrated level of efficiency the resulting democratisation could open up exciting areas for growth worldwide, favouring diverse open source solutions and providing considerable level of robustness against centralisation.
-The general nature of semantic features makes it reasonable to expect that extending these results to more complex datasets and domains should be relatively straightforward - but may require some honest engineering.
-The obvious ideas for further work include:
+4.5 Ablation study
+Most of the heavy lifting regarding pixel-intensity alignment is done by the Two Step Layer which lifts adversarial robustness from 0% to ~86%.
+The Convolutional Layer stabilizes the training and boosts adversarial robustness by further ~6%pt.
+Interestingly the number of affine semantic features can be reduced from 8 to 2 with the cost of just ~2%pt adversarial accuracy.
+Logical Layer can be replaced by a simple linear probe, however it needs a proper initialization to avoid the risk of dead neurons (which is high because of the exceedingly small amount of neurons).
+The Gaussian noise augmentation is vital for the model to experience pixel intensity variations during training (boosts adversarial robustness by ~62%pt).
+Despite being primarily designed to withstand the $L_{\infty}$ attack the model has considerable robustness against the standard $L_{2}$ AutoPGD attack (eps=1.5, eps_step=0.1) with ~89% adversarial accuracy.
+5 Discussion
+We’ve built a white box neural network that is naturally robust, reliable and inherently interpretable.
+Thanks to the notion of semantic features the construction of the model was as straightforward as writing a fully explainable procedural program and so are the inner workings of the network.
+Affine Layer is the core component of the model and the main reason for its interpretability.
+However if applied directly to the input it is easily fooled by attacks.
+It’s white box nature makes it easy to understand where this vulnerability comes from and to fix it: quick inspection of the mismatched poses as in the Figure 6 reveals that it is predominantly due to the large attack budget of $0.3$ per pixel (which is $30\%$ of the entire allowed pixel intensity range) combined with the large linear regions of the function computed by this layer (due to small amount of poses per feature).
+Human eye appears to ignore such considerate changes in pixel intensities which means that the naive linear representation of pixel brightness is not semantically adequate.
+This is exactly how (and why) the first two layers of the model were designed - to fix the inherent vulnerability of otherwise perfectly interpretable affine features.
+These layers are themselves compatible with the framework of semantic features which suggests a large untapped potential of the notion.
+The Logical Layer has been developed as an attempt to formalize the idea of an abstract hidden layer sparse activation structure and in that regard it is the most inspiring; however it might turn out to be too generic (on MNIST the properly-initialized linear probe seems to be sufficient).
+6 Further research
+Future work would have to involve more complex datasets and more nuanced semantic features.
+For example notice the analogy between affine semantic features and the SIFT [10] algorithm - the learnable pose sampling may be interpreted as finding Keypoints and their Descriptors, while previous layers (Two Step and Convolutional) are essentially performing input normalization.
+This might be a helpful analogy for designing more advanced semantic features that sample across varying scales, sharpness or illumination levels.
+There might be a concern that adding more layers to the network would reduce its interpretability.
+However the role of additional layers would be to widen the receptive field of the network.
+The Affine Layer defined earlier matches parts-at-locations of center-aligned objects.
+Adding additional layer would involve a sliding window approach (or Keypoint Detection similar to that in SIFT) to match center-aligned objects inside the window and integrate this with the approximate global location of the window; then the Global Affine Layer matches objects-at-locations which form a scene (a higher-level object).
+This would in fact be similar to applying Affine Layer to the "zoomed-out" version of the input.
+Other ideas for further work include:
   • use semantic features in self-supervised setting to obtain interpretable dimensionality reduction;
+  • study non-affine perturbations of features, e.g. projective perturbations;
+  • find adequate semantic topology for color space;
+  • define semantic features for sound (inspiration for semantic invariants could be drawn from music theory);
+  • define semantic features for text, e.g. capture semantic invariance of permutations of words;
   • study more complex logical semantic features in more elaborate scenarios;
   • figure out a way to make logical semantic features differentiable, i.e. make $\mathbf{P}$ learnable;
-  • design richer modes of spatial variation, e.g.
-hierarchical affine semantic features to capture variations of compound objects (objects consisting of affine parts are themselves affine if you "zoom out" appropriately);
-  • find adequate semantic topology for color space;
-  • integrate affine features with sliding window approach, i.e. implement weight sharing of similar features at distant locations;
-  • define semantic features for sound (inspiration for semantic invariants could be drawn from music theory);
-  • define semantic features for text, e.g. to capture semantic invariance of permutations of words;
 7 Acknowledgments
-This research has been entirely self-funded - I saved money, left my job and took the risk of following independently a depressingly underfunded and underestimated line of research.
-However it needs to be noted that I had the opportunity to do preliminary research on adversarial robustness as a half-time activity during my employment at MIM Solutions.
-Special thanks go to Piotr Sankowski, who introduced me to the issue of adversarial vulnerability, and both to Piotr Sankowski and Piotr Wygocki who - as my former employers - decided to trust me with that research as a half-time part of my contract.
+This is work was done as an independent self-funded research.
+However it needs to be noted that I had the opportunity to do preliminary investigation on adversarial robustness as a half-time activity during my employment at MIM Solutions.
+Special thanks go to Piotr Sankowski, who introduced me to the issue of adversarial vulnerability of neural networks, and both to Piotr Sankowski and Piotr Wygocki who - as my former employers - decided to trust me with that research as a half-time part of my contract.
 It was precisely thanks to this opportunity that I developed a strong intuition that modern neural network architectures are inherently flawed and need to be fundamentally re-examined.
-The story of this paper - should the results be recognized by the community - indicates a need for deep systemic and cultural changes that would empower individuals and smaller institutions to allow them to afford putting values and long-term informed risks over short-term profits.
-This is especially needed in larger public institutions like universities.
-Otherwise we might get stuck following "easy" directions which leads to a lethal combination of demoralisation and deterioration of quality.
+I thank Piotr Wodziński, Adam Szummer, Tomasz Steifer, Dmytro Mishkin, Andrzej Pacuk and members of the reddit community for the great early feedback on the first version of this paper.
 Appendix A Implementation of Two Step Layer
@@
 Appendix B Learned Affine Features with learned poses
+Appendix C Preliminary results for the entire MNIST
+Let’s review the results for a straightforward adaptation of previously defined model for the full MNIST:
+  • Affine Layer is expanded to 50 affine semantic features;
+  • Logical Layer is replaced by a linear probe with 10 output classes;
+  • Size of affine features is reduced to $14\times 14$ patches;
+  • Affine base features are equipped with learnable $(x,y)$ per-feature location;
+Such model has ~20K parameters, achieves ~98% standard and ~80% adversarial accuracy (AutoAttack) and can still be considered white box as shown in the following figures:
+Figure 7: Affine base features learned on full MNIST.
+Figure 8: Masked non-negative weights of linear probe layer - per output class.
+Figure 9: Sum of affine features weighted by positive weights of linear probe - per output class.

### v4

@@
 Abstract
-This paper introduces semantic features as a candidate conceptual framework for building inherently interpretable neural networks.
-A proof of concept model for informative subproblem of MNIST consists of 4 such layers with the total of ~5K learnable parameters.
-The model is well-motivated, inherently interpretable, requires little hyperparameter tuning and achieves human-level adversarial test accuracy - with no form of adversarial training!
+This paper introduces semantic features as a candidate conceptual framework for white box neural networks.
+The proof of concept model is well-motivated, inherently interpretable, has low parameter-count and achieves almost human-level adversarial test metrics - with no adversarial training!
 These results and the general nature of the approach warrant further research on semantic features.
@@
 1 Introduction
-Machine learning systems used in safety-critical applications such as healthcare, military, autonomous vehicles or cybersecurity need to be reliable, i.e.
-they cannot fail unpredictably.
-The black box nature of state of the art (SOTA) deep neural networks (DNN) make it exceedingly hard to give such guarantees, which limits their usability to non-critical domains.
-The recent research in Explainable AI (XAI) focuses predominantly on post-hoc attribution methods which may provide explanations but, due to the complexity of SOTA, are hard to be converted to reliable fail-safes.
-This is a burning issue because of the wide susceptibility of DNNs to distribution shifts and adversarial attacks [18, 5] meaning that models can make arbitrarily incorrect predictions and their reported high test-time metrics don’t guarantee robust real-life performance111models often rely on spurious features that leak from test to train [6]..
-The recently popular Neurosymbolic AI [17] consists of higher-level methods that focus on augmenting neural network representations (machine perception) with background structured knowledge (reasoning) but don’t address the potential failures of the former.
+The core advantages of deep neural networks (DNNs) are their architectural simplicity and automatic feature learning.
+The latter is crucial for working with unstructured data as developers don’t need to design features by hand.
+However giving away the control over features leads to black box models - DNNs tend to learn uninterpretable "shortcut" correlations [13] that leak from train to test [15], hampering alignment and out-of-distribution performance.
+In particular, this gives rise to adversarial attacks [30] - semantically negligible perturbations of data that arbitrarily change model’s predictions.
+Adversarial vulnerability is a widespread phenomenon (vision, segmentation, NLP, sound, RL, tabular data) and largely contributes to the general lack of trust in DNNs, substantially limiting their adoption in high-stakes applications such as healthcare, military, autonomous vehicles or cybersecurity.
+Conversely, the main advantage of hand-designed features is the fine-grained control over model’s performance; however, such systems quickly become infeasibly complex.
+This paper aims to address those issues by reconciling Deep Learning with feature engineering - with the help of locality engineering.
+Specifically, semantic features are introduced as a general conceptual machinery for controlled dimensionality reduction inside a neural network layer.
+Figure 1 presents the core idea behind the notion and the rigorous definition is given in Section 4.
+Implementing a semantic feature predominantly involves encoding appropriate invariants (i.e.
+semantic locality) for the desired semantic entities, e.g.
+"small affine perturbations" for shapes or "adding small $\delta\in[-\epsilon,\epsilon]$" for numbers.
+Thus, semantic features capture the core characteristic of real-world objects: having many possible states but being at exactly one state at a time.
+It turns out that for the trained proof-of-concept (PoC) DNN such parameter-sharing strategy leads to human-aligned base features - which motivates calling the PoC a white box neural network, see Section 6.
+In particular, the resulting inherent interpretability of the PoC allowed for designing a strong architectural defense111with no form of adversarial training!
+against adversarial attacks, as discussed in Section 7.
+Thus, the notion of "semantic feature" and the adversarially robust white box PoC network are the main contributions of this work.
+Figure 1: Visualisation of semantic feature $f_{\mathbf{P}\mathbf{L}}$ and its matching mechanism (SFmatch, see Section 4.2).
+The semantic feature $f_{\mathbf{P}\mathbf{L}}$ consists of a base feature $f$ and a set of its "small" modifications $\mathbf{L}(p_{i})(f)$.
+In this case, the base feature has the same dimensions as the input image and its modifications are 2D affine transformations.
+$\mathrm{SFmatch}(d,f_{\mathbf{P}\mathbf{L}})$ takes the maximum of the dot products $<d,\mathbf{L}(p_{i})(f)>$ thus identifying all $\mathbf{L}(p_{i})(f)$ with $f$.
+Affine transformations are typically parameterized by extended matrices of dimension $2\times 3$; both the base feature and the parameters $p_{i}$ of its modifications are learned.
+If $f$ and its modifications are easily understood by humans, then the neural network layer composed of such features can be considered a white box layer.
+The precise definition of semantic feature is found in Section 4.
+The paper is organised as follows.
+Section 2 investigates related work on explainability and adversarial robustness.
+Section 3 briefly describes the adopted methodology.
+Section 4 gives the rigorous definition of semantic feature and argues for the generality of the notion.
+Section 5 builds a carefully motivated PoC 4-layer white box neural network for the selected problem.
+Section 6 analyses the trained model both qualitatively and quantitatively in terms of explainability, reliability, adversarial robustness and performs an ablation study.
+Section 7 provides a short discussion of advantages and limitations of the approach.
+The paper ends with ideas for further research and acknowledgments.
+2 Related work
+2.1 Adversarial robustness
 Existing methods of improving DNN robustness are far from ideal.
 Adversarial Training [22] being one of the strongest empirical defenses against attacks [3] is slow, data-hungry [26, 18] and results in robust generalisation gap [24, 34] as models overfit to the train-time adversaries.
-Thus, the biggest challenge of empirical approaches is reliable evaluation of robustness as more and more sophisticated attacks break the previously successful defenses [3, 4].
-To address these challenges the field of certified robustness has emerged with the goal of providing theoretically certified upper bounds for adversarial error [9].
-However the exact certification is NP-Complete [20] and therefore has serious limitations in practice as it’s limited to relatively small problems and small networks and/or to probabilistic approaches.
+This highlights the biggest challenge of empirical approaches, i.e.
+reliable evaluation of robustness, as more and more sophisticated attacks break the previously successful defenses [7, 9].
+To address this issue the field of certified robustness has emerged with the goal of providing theoretically certified upper bounds for adversarial error [19].
+However the complete certification faces serious obstacles as it is NP-Complete [33] and therefore limited to small datasets and networks, and/or to probabilistic approaches.
 Another shortcoming is that such guarantees can only be given for a precisely defined threat model which is usually assumed to be a $L_{p}$-bounded perturbation.
@@
 for different $p\in\{0,1,2,\ldots,\infty\}$ and often require training the model specifically towards chosen $L_{p}$-certification.
-The highlighted issues are relevant for datasets as simple as MNIST [16] or even some of its binary subsets [19] - despite being "toy" and almost linearly separable problems, they still remain elusive in terms of robustness, rising further questions about the true generalisation capabilities of SOTA as no neural network model can be honestly considered reliable and adequate333in terms of internal complexity; SOTA models define inappropriately complicated functions.
-for this allegedly "simple" problem.
-1.1 Proposed approach
-This paper proposes an intermediate approach to DNN reliability and introduces a candidate theoretical framework for building white box neural networks.
-Building and analysing a white box DNN would be similar to writing and debugging a procedural program - it doesn’t have to come with theoretical certificates but needs to be easily understandable and internalizable by the programmer.
-A good software has to be designed well enough to account for reasonable failure scenarios, handle foreseeable exceptions and, should an unforseen bug happen, be amenable for fixes.
-If the design and inner mechanics of a DNN are sufficiently clear then working with such model wouldn’t be very different from maintaining a standard application.
-Current SOTA models are large, unwieldy black boxes that require prohibitive amounts of resources; yet animal brains apparently learn rich and robust patterns in mostly unsupervised sensory input - even brains as simple as drosophila brain with ~135K neurons!
-This scale seems easily maintainable with proper abstractions.
-After all, the core algorithms used today are mostly unchanged since ~30 years!
-What if there really exists a better way?
-To find it we must start with the basics and take the agile approach.
-Focusing on the simplest relevant problem cuts off the obfuscating details and speeds up the experimentation phase.
-With that in mind the Minimum Viable Dataset (MVD) for this paper is chosen to be the binary subset of MNIST consisting of images of "3" and "5" - as this is one of the most often confused pairs under adversaries [19] and therefore might be considered harder than the entire MNIST in terms of average adversarial accuracy444the preliminary results on the full MNIST are discussed in Appendix C..
-1.2 The core idea
-This paper introduces semantic features as a general idea for sharing weights inside a neural network layer.
-The rigorous definition will be given in the next section.
-The concept is similar to that of "inverse rendering" in Capsule Networks [14] where features have many possible states and the best-matching state has to be found.
-Identifying different states by the subsequent layers gives rise to controlled dimensionality reduction.
-Thus semantic features aim to capture the core characteristic of any semantic entity - having many possible states but being at exactly one state at a time.
-This is in fact a pretty strong regularization.
-As shown in this paper, choosing appropriate layers of semantic features for the MVD results in what can be considered as a white box neural network.
-The remainder of the paper is structured as follows:
-  • Section 2. gives a rigorous definition of semantic feature and argues for the generality of the notion;
-  • Section 3. builds a carefully motivated PoC 4-layer white box neural network for selected MVD;
-  • Section 4. analyses the trained model both qualitatively and quantitatively in terms of explainability, reliability, adversarial robustness and ablation;
-  • Section 5. provides a short discussion of the results;
-The paper ends with ideas for further research and acknowledgments.
+2.2 Explainability
+Recent work indicates a link between adversarial robustness and inherent interpretability as adversarially trained models tend to enjoy Perceptually-Aligned Gradients (PAG) - taking gradient steps along the input to maximise the activation of a selected neuron in the representation space often yields visually aligned features [10, 17].
+There is also some evidence for the opposite implication that PAG implies certain form of adversarial robustness [12, 29].
+However Section 6.4 of this paper shows that the latter implication does not hold - model can be inherently interpretable but not robust.
+Susceptibility to adversarial attacks is arguably the main drawback of known approaches to Explainable AI (XAI).
+One would reasonably expect explainable models to rely exclusively on human-aligned features instead of spurious "shortcut" signals [13] and therefore be robust to adversaries.
+However the post-hoc attribution methods, which aim to explain the decisions of already-trained models, are fragile [1] and themselves prone to attacks [14].
+On the other hand the inherently interpretable solutions such as Prototypical Parts [20, 8], Bag-of-Features [5], Self-Explaining Networks [2], SITE [32], Recurrent Attention [11], Neurosymbolic AI [28] or Capsule Networks [25] all employ various classic black box encoder networks and thus introduce the adversarial vulnerability into their cores.
+Overall the field of adversarially robust explanations (AdvXAI) is fairly unexplored; for more information on this emerging research domain see the recent survey [4].
+3 Methodology
+This paper focuses on the simplest relevant problem (Minimum Viable Dataset, MVD) to cut off the obfuscating details and get into the core of what makes neural networks fragile and uninterpretable.
+Despite much effort, DNNs remain susceptible to adversarial attacks even on MNIST [27] and therefore cannot be considered human-aligned, see Figure 2.
+That being said, the MVD for this paper is chosen to be the binary subset of MNIST consisting of images of "3" and "5" - as this is one of the most often confused pairs under adversaries [31] and might be considered harder than the entire MNIST in terms of average adversarial accuracy.
+For completeness, the preliminary results on the full MNIST are discussed in Appendix C.
+The paper provides both qualitative and quantitative results.
+The qualitative analysis focuses on visualising the internal distribution of weights inside layers and how features are matched against the input.
+For the quantitative results a strong adversarial regime is employed besides computing the standard clean test metrics (computing the clean metrics alone is considered a flawed approach for the reasons discussed earlier).
+Figure 2: A typical DNN for MNIST can be arbitrarily fooled by adding semantically negligible noise.
 4 Semantic features
 In machine learning inputs are represented as tensors.
-Yet the standard Euclidean topology in tensor space usually fails to account for small domain-specific variations of features.
+Yet, the standard Euclidean topology in tensor space usually fails to account for small domain-specific variations of features.
 For example shifting an image by 1 pixel is semantically negligible but usually results in a distant output in $L_{2}$ metric.
@@
 A semantic feature $f_{\mathbf{P}\mathbf{L}}$ of dimension $s\in\mathbb{N}$ is a tuple $(f,\mathbf{P},\mathbf{L})$ where:
-  • $f$ base, $f\in\mathbb{R}^{s}$,
+  • $f$ base feature, $f\in\mathbb{R}^{s}$,
   • $\mathbf{P}$ parameter set, $\mathbf{P}\subset\mathbb{R}^{r}$ for some $r\in\mathbb{N}$
@@
 Additionally the feature $(f,\mathbf{P},\mathbf{L})$ is called differentiable iff $\mathbf{L}$ is differentiable (on some neighbourhood of $\mathbf{P}$).
-Intuitively a semantic feature consists of a base vector $f$ and a set of it’s "small" variations $f_{p}=\mathbf{L}(p)(f)$ for all $p\in\mathbf{P}$; these variations may be called poses after [14].
+Intuitively, semantic feature consists of a base vector $f$ and a set of it’s "small" variations $f_{p}=\mathbf{L}(p)(f)$ for all $p\in\mathbf{P}$; these variations may be called poses after [25].
 Thus $(f,\mathbf{P},\mathbf{L})$ is locality-sensitive in the adequate topology capturing semantics of the domain.
 The differentiability of $\mathbf{L}(p)$ and of $\mathbf{L}$ itself allow for gradient updates of $f$ and $\mathbf{P}$ respectively.
-The intention is to learn both $f$ and $\mathbf{P}$ while defining $\mathbf{L}$ explicitly as an inductive bias555by the ”no free lunch” theorem a general theory of learning systems has to adequately account for the inductive bias.
+The intention is to learn both $f$ and $\mathbf{P}$ while defining $\mathbf{L}$ explicitly as an inductive bias333by the ”no free lunch” theorem a general theory of learning systems has to adequately account for the inductive bias.
 for the given modality.
@@
     • $\mathbf{P}$ set of $2\times 2$ rotation matrices
-    • $\mathbf{L}$ function that maps a rotation matrix to the respective rotation of a 2D kernel666can be made differentiable as in [12, 7].
+    • $\mathbf{L}$ function that maps a rotation matrix to the respective rotation of a 2D kernel444can be made differentiable as in [23, 16].
   3. 3.
@@
     • $f$ vector representing a small 2D spatial feature (e.g. image patch containing a short line)
-    • $\mathbf{P}$ set of $2\times 3$ invertible augmented matrices777preferably close to the identity matrix.
-    • $\mathbf{L}$ function that maps a $2\times 3$ matrix to the respective 2D affine transformation of images6
+    • $\mathbf{P}$ set of $2\times 3$ invertible augmented matrices555preferably close to the identity matrix.
+    • $\mathbf{L}$ function that maps a $2\times 3$ matrix to the respective 2D affine transformation of images4
   4. 4.
@@
 If $\mathbf{L}$ captures domain topology adequately then we may think of matching $f_{\mathbf{P}\mathbf{L}}$ as a form of local inhibition along $\mathbf{L}_{\mathbf{P}}(f)=\{L(p)(f),p\in P\}$ and in this sense semantic feature defines a XOR gate over $\mathbf{L}_{\mathbf{P}}(f)$.
-On the other hand every $f_{p}\in\mathbf{L}_{\mathbf{P}}(f)$ can be viewed as a conjunction (AND gate) of its non-zero coordinates888i.e.
+On the other hand every $f_{p}\in\mathbf{L}_{\mathbf{P}}(f)$ can be viewed as a conjunction (AND gate) of its non-zero coordinates666i.e.
 $f_{p}$ is a conjunction of lower-level features..
@@
   • for real-valued $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=\mathrm{int}(d==g)$ where $d,g\in\mathbb{R}$
-  • for convolutional $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=\max_{0<=i<c_{out}}(d*\frac{g^{i}}{||g^{i}||_{2}})$ where $g$ is a 2D convolutional kernel of shape $(c_{out},c_{in},k,k)$, $g^{i}$ is the $i$-th convolutional filter of $g$, $d$ is an image patch of shape $(c_{in},k,k)$ 999alternatively we could treat every $g^{i}$ as a separate semantic feature but will we stick to the max option for simplicity.
+  • for convolutional777an alternative definition would treat every $g^{i}$ as a separate semantic feature; we stick to the max option for simplicity.
+$f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=\max_{0<=i<c_{out}}(d*\frac{g^{i}}{||g^{i}||_{2}})$ where $g$ is a 2D convolutional kernel of shape $(c_{out},c_{in},k,k)$, $g^{i}$ is the $i$-th convolutional filter of $g$, $d$ is an image patch of shape $(c_{in},k,k)$
   • for affine $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=d\cdot\frac{g}{||g||_{2}}$ where $g$ is a 2D image of shape $(c_{in},k,k)$ and $d$ is a 2D image patch of the same shape
@@
 Note that a single neural network layer will consist of many parallel semantic features of the same kind.
-The network consists of the following 4 layers stacked sequentially. Note that for simplicity we don’t add bias to any of those layers. Te resulting model has around $5$K parameters.
+The network consists of the following 4 layers stacked sequentially.
+For simplicity we don’t add bias to any of those layers.
+Te resulting model has around $5$K parameters.
+The architecture is visualized in the Figure 3.
+Figure 3: Architecture of the PoC white box neural network.
+The first two layers consist of semantic features that operate per-pixel - the first layer takes into account only the pixel’s value, while the second examines its $5\times 5$ neighborhood to determine if the pixel lays on a "bright line".
+The first two layers retain the shape of the input image.
+The third layer comprises of 8 affine $f_{\mathbf{P}\mathbf{L}}$ as visualized in Figure 1.
+The final layer consists of two logical $f_{\mathbf{P}\mathbf{L}}$: intuitively, the first one checks whether at least one affine $f_{\mathbf{P}\mathbf{L}}$ corresponding to "3" is active and none of the affine $f_{\mathbf{P}\mathbf{L}}$ corresponding to "5" are active; the second logical $f_{\mathbf{P}\mathbf{L}}$ works in the opposite way.
+Section 5 describes the architecture in more detail.
 5.1 Two Step Layer
@@
 Its basic building block is a pixel $x\in[0,1]$.
-Despite being allowed to assume any value between 00 and $1$ it is actually (semantically) a ternary object - it’s either ON, OFF or MEH101010the MEH state is an undecided state in-between..
+Despite being allowed to assume any value between 00 and $1$ it is actually (semantically) a ternary object - it’s either ON, OFF or MEH888the MEH state is an undecided state in-between..
 Therefore the semantic space should squash the $[0,1]$ interval into those 3 values.
@@
 We initialize the layer setting init_scales_div=10.
-3.2 Convolutional Layer
+5.2 Convolutional Semantic Layer
 A pixel-wise function is not enough to classify pixel as ON or OFF - the semantics of our MVD require a pixel to exist in a sufficiently bright region to be considered as ON.
@@
 The layer performs $\mathrm{SFmatch}(d(x,y),f_{\mathbf{P}\mathbf{L}})$ with every pixel $(x,y)$ of image $d$.
-This means that it matches the two rotated filters $g^{0}$ and $g^{1}$ with $(x,y)$ and takes the maximum across all the matches111111a single number, not a pair of numbers, as in the chosen definition of match for convolutional $f_{\mathbf{P}\mathbf{L}}$..
-Intuitively this layer checks if the pixel has a sufficiently bright neighbourhood.
-The layer is followed by ReLU activation. We initialize $g^{0}$ and $g^{1}$ as the identity kernel of shape $(1,1,5,5)$ plus Gaussian noise $\sim\mathcal{N}(0,0.1)$.
+This means that it matches the two rotated filters $g^{0}$ and $g^{1}$ with the $5\times 5$ neighbourhood of $(x,y)$ and takes the maximum across all the matches999a single number, not a pair of numbers, as in the chosen definition of match for convolutional $f_{\mathbf{P}\mathbf{L}}$..
+Intuitively, this layer checks if pixels have adequately structured neighbourhoods.
+The layer is followed by ReLU activation. We initialize $g^{0}$ and $g^{1}$ each as the identity kernel of shape $(1,1,5,5)$ plus Gaussian noise $\sim\mathcal{N}(0,0.1)$.
 5.3 Affine Layer
@@
 In short they are shapes at locations.
-The semantic identity of shape is not affected by small121212close to the identity augmented matrix affine transformations.
+The semantic identity of shape is not affected by small101010close to the identity augmented matrix.
+affine transformations.
 The next shape-extracting layer will therefore consist of 8 affine semantic features $(f,\mathbf{P},\mathbf{L})$ defined as follows:
@@
 For simplicity every $f_{\mathbf{P}\mathbf{L}}$ in this layer is zero-padded evenly on every side and identified with the resulting $28\times 28$ image.
-Intuitively this matches the image against a localized shape in a way that is robust to small affine perturbations of the shape.
+Intuitively, this matches the image against a localized shape in a way that is robust to small affine perturbations of the shape.
 The layer is followed by ReLU activation.
@@
 After Affine Layer has extracted predictive shapes it remains to encode logical claims such as "number 3 is A or B and not C and not D" etc.
-Since the previous layer is learnable we can enforce a preferable structure on the input neurons to the Logical Layer.
+Since the previous layer is learnable, we can enforce a preferable structure on the input neurons to the Logical Layer.
 This layer will consist of 2 logical $f_{\mathbf{P}\mathbf{L}}$ (one for every label) defined as follows:
@@
 The model was trained to optimize the CrossEntropyLoss for 15 epochs with Adam optimizer, learning rate $3e{-3}$, weight decay $3e{-6}$.
-The only augmentation was Gaussian noise $\sim\mathcal{N}(0.05,0.25)$ applied with $0.7$ probability131313mean is slightly positive because of the sparsity of data - subtracting pixel value is on average more harmful than adding it..
+The only augmentation was Gaussian noise $\sim\mathcal{N}(0.05,0.25)$ applied with $0.7$ probability111111mean is slightly positive because of the sparsity of data - subtracting pixel value is on average more harmful than adding it..
 Those parameters were chosen "by hand" after few rounds of experimentation.
@@
 Learning curve
-Figure 1: (left) Reliability test-time curve for the classic 40-step PGD Attack. (right) Learning curve.
+Figure 4: (left) Reliability test-time curve (see Section 6.2 for details). (right) Learning curve. Both test and validation metrics presented here are computed under the classic 40-step PGD Attack.
 6.2 Quantitative results
-We don’t bother measuring the clean test metrics as it is a flawed approach that tells little about the true generalisation capability of the model.
-Every metric reported here is computed under strong adversarial regime.
-The model achieves ~92% accuracy under AutoAttack [4] (with default parameters: norm=’inf’, eps=0.3, step_size=0.1).
-The metrics are almost identical for the much faster 40-step PGD Attack.
-Samples are usually misclassified with low confidence and therefore we compute a reliability curve - a precision-recall curve where the positive class is defined as the set of correctly classified samples (this means that the precision for 100% recall is exactly equal to the accuracy).
-It turns out that for 80% recall model achieves human-level 98% adversarial precision (Figure 1).
-This means that in practice human intervention would be required only for the easily detected 20% of adversarial samples to achieve human level reliability of the entire system (and most real-life samples are not adversarial).
+The model achieves ~92% adversarial test accuracy under AutoAttack [9] (with the default parameters: norm=’inf’, eps=0.3, step_size=0.1).
+The results are almost identical for the much faster 40-step PGD Attack with the same parameters.
+Adversarial samples are usually misclassified with low confidence and therefore the reliability curve is computed for the PGD-attacked test images, i.e.
+a precision-recall curve where the positive class is defined as the set of correctly classified samples (this means that the precision for 100% recall is exactly equal to the accuracy).
+It turns out that for 80% adversarial recall model achieves human-level 98% adversarial precision (Figure 4).
+This means that in practice human intervention would be required only for the easily detected ~20% of adversarial samples to achieve human-level reliability of the entire system (and most real-life samples are not adversarial).
+The clean test accuracy is ~99.5%.
 6.3 Layer inspection
-Lets inspect the trained model layers.
 6.3.1 Two Step Layer
-Figure A shows the initial and learned shapes of the Two Step function.
+Figure 5 shows the initial and learned shapes of the Two Step function.
 The function smoothly thresholds input at $0.365$ and then at $0.556$ which means that the interval $[0,0.365]$ is treated as OFF, $[0.365,0.556]$ as MEH and $[0.556,1]$ as ON.
@@
 Figure 5: Two Step layer - initial and learned.
-4.3.2 Convolutional Layer
-Quick inspection of the Figure 3 shows that the Convolutional Layer has indeed learned meaningful filters that check if the central pixel lays inside an adequately structured region - intuitively the matching value corresponds to the degree the model considers the pixel as being ON.
+The visible smooth thresholding essentially means that the layer has learned the permissible perturbations (i.e.
+intervals) of real semantic features corresponding to the state of being OFF (number -1) and ON (number 1).
+6.3.2 Convolutional Semantic Layer
+Quick inspection of the Figure 6 shows that the Convolutional Semantic Layer has indeed learned meaningful filters that check if the central pixel lays inside an adequately structured region.
 Figure 6: Learned convolutional kernels in 32 predefined poses.
+Convolutional Semantic Layer acts pixel-wise and matches (convolves) these rotated kernels against the $5\times 5$ neighbourhood of a given pixel, returning the maximum matching value.
+Intuitively, this corresponds to the degree of pixel laying on a "bright line".
 6.3.3 Affine Layer
-Figure 4 shows the learned affine base features. Figure B in the Appendix B shows learned features in 32 learned feature-specific poses. It’s hard to imagine substantially more meaningful features.
-Figure 4: Learned affine base features.
+Figure 7 shows the learned affine base features.
+Appendix B shows the 8 learned affine features in all of the 32 learned feature-specific poses.
+It’s hard to imagine substantially more meaningful features.
+Figure 7: Learned affine base features. The weight-sharing strategy imposed by the affine locality function results in easily interpretable features.
 6.3.4 Logical Layer
@@
 For $f_{\mathbf{P}\mathbf{L}}^{1}$ its the other way round.
-4.4 Decision boundary
-Figure 5 shows the minimum perturbations required to change model’s predictions under Boundary Attack [2] for sample test images.
-The added perturbations are easily spotted and understood by humans.
-Note that by definition model is extremely unconfident on those samples and in practice the perturbations required to fool the entire system would have to be much larger.
-Figure 5: Boundary Attack: minimal perturbations that flip model prediction (model confidence at around 50%).
-Figure 6: Visualization of test data as "seen" by the first 3 consecutive layers.
-Figure 6 shows how the data is "seen" by the first 3 layers. The consecutive rows and labels denote respectively:
-  1. 1.
-    adversarial input and predicted class probabilities
-  2. 2.
-    output of Two Step Layer
-  3. 3.
-    output of Convolutional Layer
-  4. 4.
-    top pose of affine features predicting label "3" (with the match value)141414these rows are actually overlaid with the middle row to make it easier to see what has been matched.
-  5. 5.
-    top pose of affine features predicting label "5" (with the match value)14
-Note that in Figure 6 there are 3 misclassified examples and for every one of them it’s easy to understand why the model made the mistake.
+Figure 8: Minimally perturbated images under Boundary Attack [6] to change model’s predictions - for sample test images.
+The perturbations are often meaningful to humans and introduce semantic ambiguity.
+Note that by definition of the Boundary Attack the model is extremely unconfident on those samples (model confidence at ~50%) and in practice the perturbations required to fool the entire system would have to be much larger.
+Figure 9: Visualization of test data as "seen" by the first 3 layers.
+The consecutive rows and labels denote respectively: adversarial input and predicted class probabilities; output of Two Step Layer; output of Convolutional Semantic Layer; top pose of affine features predicting label "3" (with the match value above); same for label "5".
+The last two rows are overlaid with the middle row to make it easier to see what has been matched.
+Note that there are 3 misclassified examples (first, 8th and last) and for every one of them it’s easy to understand why the model made the mistake: for the first one the model recognizes the middle part of "3" as the top part of "5"; the "5" in the 8th image has specific curve at the top which resembles the top part of "3"; the "3" in the last example has very thin top part which gets removed by the first 2 layers thus creating an ambiguity.
 6.4 Ablation study
 Most of the heavy lifting regarding pixel-intensity alignment is done by the Two Step Layer which lifts adversarial robustness from 0% to ~86%.
-The Convolutional Layer stabilizes the training and boosts adversarial robustness by further ~6%pt.
+The Convolutional Semantic Layer stabilizes the training and boosts adversarial robustness by further ~6%pt.
 Interestingly the number of affine semantic features can be reduced from 8 to 2 with the cost of just ~2%pt adversarial accuracy.
@@
 Thanks to the notion of semantic features the construction of the model was as straightforward as writing a fully explainable procedural program and so are the inner workings of the network.
+7.1 Defending against adversaries
 Affine Layer is the core component of the model and the main reason for its interpretability.
 However if applied directly to the input it is easily fooled by attacks.
-It’s white box nature makes it easy to understand where this vulnerability comes from and to fix it: quick inspection of the mismatched poses as in the Figure 6 reveals that it is predominantly due to the large attack budget of $0.3$ per pixel (which is $30\%$ of the entire allowed pixel intensity range) combined with the large linear regions of the function computed by this layer (due to small amount of poses per feature).
+Its white box nature makes it easy to understand where this vulnerability comes from and to fix it: quick inspection of the mismatched poses as in the Figure 9 reveals that it is predominantly due to the large attack budget of $0.3$ per pixel (which is $30\%$ of the entire allowed pixel intensity range) combined with the large linear regions of the function computed by this layer (due to small amount of poses per feature).
 Human eye appears to ignore such considerate changes in pixel intensities which means that the naive linear representation of pixel brightness is not semantically adequate.
 This is exactly how (and why) the first two layers of the model were designed - to fix the inherent vulnerability of otherwise perfectly interpretable affine features.
-These layers are themselves compatible with the framework of semantic features which suggests a large untapped potential of the notion.
-The Logical Layer has been developed as an attempt to formalize the idea of an abstract hidden layer sparse activation structure and in that regard it is the most inspiring; however it might turn out to be too generic (on MNIST the properly-initialized linear probe seems to be sufficient).
+These layers are themselves compatible with the framework of semantic features which is an early indication of a large potential of the notion.
+7.2 Limitations
+Semantic features require locality engineering as they rely on the adequate domain-specific definition of the locality function.
+Therefore they won’t be applicable in contexts where such locality is hard to define.
+On the other hand it opens up interesting areas for research.
+The Logical Layer has been developed as an attempt to formalize the idea of how the generic hidden layer activation pattern should look like; however it might turn out to be prone to the "no free lunch" theorem (e.g.
+on MNIST the properly-initialized linear probe seems to be sufficient).
+Despite the generality of the notion and the ideas given in Section 8 there is a risk that the white box qualities of semantic features will be difficult to achieve for harder datasets.
+The preliminary results on CIFAR10 seem promising but also reveal the need for careful engineering of the locality function, see Section 8.
 8 Further research
 Future work would have to involve more complex datasets and more nuanced semantic features.
-For example notice the analogy between affine semantic features and the SIFT [10] algorithm - the learnable pose sampling may be interpreted as finding Keypoints and their Descriptors, while previous layers (Two Step and Convolutional) are essentially performing input normalization.
+For example notice the analogy between affine semantic features and the SIFT [21] algorithm - the learnable pose sampling may be interpreted as finding Keypoints and their Descriptors, while previous layers (Two Step and Convolutional Semantic) are essentially performing input normalization.
 This might be a helpful analogy for designing more advanced semantic features that sample across varying scales, sharpness or illumination levels.
@@
 However the role of additional layers would be to widen the receptive field of the network.
-The Affine Layer defined earlier matches parts-at-locations of center-aligned objects.
+The Affine Layer defined earlier matches shapes-at-locations of center-aligned objects.
 Adding additional layer would involve a sliding window approach (or Keypoint Detection similar to that in SIFT) to match center-aligned objects inside the window and integrate this with the approximate global location of the window; then the Global Affine Layer matches objects-at-locations which form a scene (a higher-level object).
 This would in fact be similar to applying Affine Layer to the "zoomed-out" version of the input.
-Other ideas for further work include:
-  • use semantic features in self-supervised setting to obtain interpretable dimensionality reduction;
-  • study non-affine perturbations of features, e.g. projective perturbations;
-  • find adequate semantic topology for color space;
-  • define semantic features for sound (inspiration for semantic invariants could be drawn from music theory);
-  • define semantic features for text, e.g. capture semantic invariance of permutations of words;
-  • study more complex logical semantic features in more elaborate scenarios;
-  • figure out a way to make logical semantic features differentiable, i.e. make $\mathbf{P}$ learnable;
+The ablations in Section 6.4 suggest that adversarial vulnerability of DNNs might be mostly due to the fragility of the initial layers.
+Therefore another direction forward would be to re-examine the known approaches to XAI and try to "bug fix" their encoder networks in similar fashion.
+Other ideas for further work on semantic features include:
+  • use semantic features in self-supervised setting to obtain interpretable dimensionality reduction
+  • study non-affine perturbations of features, e.g. projective perturbations
+  • find adequate semantic topology for color space
+  • semantic features for sound (inspiration for semantic invariants could be drawn from the music theory)
+  • semantic features for text, e.g. semantic invariants of permutations of words/sentences capturing thoughts
+  • study more complex logical semantic features in more elaborate scenarios
+  • figure out a way to make logical semantic features differentiable, i.e. make $\mathbf{P}$ learnable
 9 Acknowledgments
-This is work was done as an independent self-funded research.
+This is work has been done as an independent self-funded research.
 However it needs to be noted that I had the opportunity to do preliminary investigation on adversarial robustness as a half-time activity during my employment at MIM Solutions.
@@
 It was precisely thanks to this opportunity that I developed a strong intuition that modern neural network architectures are inherently flawed and need to be fundamentally re-examined.
-I thank Piotr Wodziński, Adam Szummer, Tomasz Steifer, Dmytro Mishkin, Andrzej Pacuk and members of the reddit community for the great early feedback on the first version of this paper.
+I thank Piotr Wodziński, Adam Szummer, Tomasz Steifer, Dmytro Mishkin, Andrzej Pacuk, Bartosz Zieliński, Hubert Baniecki and members of the reddit community for the great feedback on the initial versions of this paper.
 Appendix A Implementation of Two Step Layer
@@
 Such model has ~20K parameters, achieves ~98% standard and ~80% adversarial accuracy (AutoAttack) and can still be considered white box as shown in the following figures:
-Figure 7: Affine base features learned on full MNIST.
-Figure 8: Masked non-negative weights of linear probe layer - per output class.
-Figure 9: Sum of affine features weighted by positive weights of linear probe - per output class.
+Figure 10: Affine base features learned on full MNIST. Due to the learnable feature-specific $(x,y)$-location these can be considered shapes-at-locations.
+Figure 11: Masked positive weights of linear probe - per output class (here the dark-blue squares are zeros).
+There are only few dark red neurons per class which means that the digit representation can be considered sparse.
+Figure 12: Sum of base affine features weighted by positive weights of linear probe - per output class.
+The similarity to actual digits means that the model decomposes the digits into human-aligned shapes-at-locations.

### v5

@@
 Abstract
-This paper introduces semantic features as a candidate conceptual framework for white box neural networks.
-The proof of concept model is well-motivated, inherently interpretable, has low parameter-count and achieves almost human-level adversarial test metrics - with no adversarial training!
+Deep neural networks learn fragile "shortcut" features, rendering them difficult to interpret (black box) and vulnerable to adversarial attacks.
+This paper proposes semantic features as a general architectural solution to this problem.
+The main idea is to make features locality-sensitive in the adequate semantic topology of the domain, thus introducing a strong regularization.
+The proof of concept network is lightweight, inherently interpretable and achieves almost human-level adversarial test metrics - with no adversarial training!
 These results and the general nature of the approach warrant further research on semantic features.
@@
 1 Introduction
-The core advantages of deep neural networks (DNNs) are their architectural simplicity and automatic feature learning.
+The main advantages of deep neural networks (DNNs) are their architectural simplicity and automatic feature learning.
 The latter is crucial for working with unstructured data as developers don’t need to design features by hand.
-However giving away the control over features leads to black box models - DNNs tend to learn uninterpretable "shortcut" correlations [13] that leak from train to test [15], hampering alignment and out-of-distribution performance.
+However, giving away the control over features leads to black box models - DNNs tend to learn hardly interpretable "shortcut" correlations [17] that leak from train to test [20], hampering alignment and out-of-distribution performance.
 In particular, this gives rise to adversarial attacks [35] - semantically negligible perturbations of data that arbitrarily change model’s predictions.
-Adversarial vulnerability is a widespread phenomenon (vision, segmentation, NLP, sound, RL, tabular data) and largely contributes to the general lack of trust in DNNs, substantially limiting their adoption in high-stakes applications such as healthcare, military, autonomous vehicles or cybersecurity.
+Adversarial vulnerability is a widespread phenomenon (vision [35], segmentation/detection [39], speech recognition [9], tabular data [10], RL [19], NLP [41]) and largely contributes to the general lack of trust in DNNs, substantially limiting their adoption in high-stakes applications such as healthcare, military, autonomous vehicles or cybersecurity.
 Conversely, the main advantage of hand-designed features is the fine-grained control over model’s performance; however, such systems quickly become infeasibly complex.
@@
 In this case, the base feature has the same dimensions as the input image and its modifications are 2D affine transformations.
-$\mathrm{SFmatch}(d,f_{\mathbf{P}\mathbf{L}})$ takes the maximum of the dot products $<d,\mathbf{L}(p_{i})(f)>$ thus identifying all $\mathbf{L}(p_{i})(f)$ with $f$.
+$\mathrm{SFmatch}(d,f_{\mathbf{P}\mathbf{L}})$ takes the maximum of the scalar products $d\cdot\mathbf{L}(p_{i})(f)$ thus identifying all $\mathbf{L}(p_{i})(f)$.
 Affine transformations are typically parameterized by extended matrices of dimension $2\times 3$; both the base feature and the parameters $p_{i}$ of its modifications are learned.
@@
 The methods, results and theoretical obstacles can be vastly different222for example SOTA certified robustness for MNIST is ~94% for $L_{\infty},eps=0.3$ and only ~73% for $L_{2},eps=1.58$, see https://sokcertifiedrobustness.github.io/leaderboard/.
-for different $p\in\{0,1,2,\ldots,\infty\}$ and often require training the model specifically towards chosen $L_{p}$-certification.
+for different $p\in\{0,1,2,\ldots,\infty\}$ and often require training the model specifically towards the chosen $L_{p}$-certification.
 2.2 Explainability
@@
 There is also some evidence for the opposite implication that PAG implies certain form of adversarial robustness [15, 34].
-However Section 6.4 of this paper shows that the latter implication does not hold - model can be inherently interpretable but not robust.
+However, Section 6.4 of this paper shows that the latter implication does not hold - model can be inherently interpretable but not robust.
 Susceptibility to adversarial attacks is arguably the main drawback of known approaches to Explainable AI (XAI).
 One would reasonably expect explainable models to rely exclusively on human-aligned features instead of spurious "shortcut" signals [17] and therefore be robust to adversaries.
-However the post-hoc attribution methods, which aim to explain the decisions of already-trained models, are fragile [1] and themselves prone to attacks [14].
-On the other hand the inherently interpretable solutions such as Prototypical Parts [20, 8], Bag-of-Features [5], Self-Explaining Networks [2], SITE [32], Recurrent Attention [11], Neurosymbolic AI [28] or Capsule Networks [25] all employ various classic black box encoder networks and thus introduce the adversarial vulnerability into their cores.
+However, the post-hoc attribution methods, which aim to explain the decisions of already-trained models, are fragile [1] and themselves prone to attacks [18].
+On the other hand, the inherently interpretable solutions such as Prototypical Parts [25, 11], Bag-of-Features [5], Self-Explaining Networks [2], SITE [37], Recurrent Attention [14], Neurosymbolic AI [33] or Capsule Networks [30] all employ various classic black box encoder networks and thus introduce the adversarial vulnerability into their cores.
 Overall the field of adversarially robust explanations (AdvXAI) is fairly unexplored; for more information on this emerging research domain see the recent survey [4].
+2.3 Geometric Deep Learning
+The idea behind semantic features resembles the framework of Geometric Deep Learning [7] or the more recent Categorical Deep Learning [16] in that semantic features are invariant to small deformations (local symmetries).
+However, semantic features are more practical as they focus on the local invariance and don’t require defining a formal structure on the input space; instead, they learn to sample the semantic neighbourhood of a base feature (which is also learned).
+Thus, the only effort required from the developer is to implement sufficient sampling mechanism, which doesn’t need to be formal or exhaustive, e.g.
+small affine perturbations seem to be good enough for MNIST features (see Section 6.3.3), despite the fact that not every possible local perturbation of these features is affine.
+Additionally, the mentioned frameworks are high-level approaches that don’t address explicitly the core issue of adversarial vulnerability.
 3 Methodology
 This paper focuses on the simplest relevant problem (Minimum Viable Dataset, MVD) to cut off the obfuscating details and get into the core of what makes neural networks fragile and uninterpretable.
-Despite much effort, DNNs remain susceptible to adversarial attacks even on MNIST [27] and therefore cannot be considered human-aligned, see Figure 2.
+Despite much effort, DNNs remain susceptible to adversarial attacks even on MNIST [32] and therefore cannot be considered human-aligned even for this simple dataset, see Figure 2.
 That being said, the MVD for this paper is chosen to be the binary subset of MNIST consisting of images of "3" and "5" - as this is one of the most often confused pairs under adversaries [36] and might be considered harder than the entire MNIST in terms of average adversarial accuracy.
@@
   • $\mathbf{P}$ parameter set, $\mathbf{P}\subset\mathbb{R}^{r}$ for some $r\in\mathbb{N}$
-  • $\mathbf{L}$ locality function : $\mathbf{P}\rightarrow\textrm{Aut}(\mathbb{R}^{s},\mathbb{R}^{s})$ where Aut is the family of differentiable automorphisms of $\mathbb{R}^{s}$
+  • $\mathbf{L}$ locality function : $\mathbf{P}\rightarrow\textrm{Diff}(\mathbb{R}^{s})$ where Diff is the family333informally, $\mathbf{L}(\mathbf{P})$ should consist of ”small” diffeomorphisms.
+of differentiable automorphisms of $\mathbb{R}^{s}$
 Additionally the feature $(f,\mathbf{P},\mathbf{L})$ is called differentiable iff $\mathbf{L}$ is differentiable (on some neighbourhood of $\mathbf{P}$).
@@
 The differentiability of $\mathbf{L}(p)$ and of $\mathbf{L}$ itself allow for gradient updates of $f$ and $\mathbf{P}$ respectively.
-The intention is to learn both $f$ and $\mathbf{P}$ while defining $\mathbf{L}$ explicitly as an inductive bias333by the ”no free lunch” theorem a general theory of learning systems has to adequately account for the inductive bias.
+The intention is to learn both $f$ and $\mathbf{P}$ while defining $\mathbf{L}$ explicitly as an inductive bias444by the ”no free lunch” theorem a general theory of learning systems has to adequately account for the inductive bias.
 for the given modality.
@@
     • $\mathbf{P}$ set of $2\times 2$ rotation matrices
-    • $\mathbf{L}$ function that maps a rotation matrix to the respective rotation of a 2D kernel444can be made differentiable as in [23, 16].
+    • $\mathbf{L}$ function that maps a rotation matrix to the respective rotation of a 2D kernel555can be made differentiable as in [28, 21].
   3. 3.
@@
     • $f$ vector representing a small 2D spatial feature (e.g. image patch containing a short line)
-    • $\mathbf{P}$ set of $2\times 3$ invertible augmented matrices555preferably close to the identity matrix.
-    • $\mathbf{L}$ function that maps a $2\times 3$ matrix to the respective 2D affine transformation of images4
+    • $\mathbf{P}$ set of $2\times 3$ invertible augmented matrices666preferably close to the identity matrix.
+    • $\mathbf{L}$ function that maps a $2\times 3$ matrix to the respective 2D affine transformation of images5
   4. 4.
@@
 The role of semantic features is to be matched against the input to uncover the general structure of the data.
-Suppose that we have already defined a function match: $(\mathbb{R}^{c},\mathbb{R}^{s})\rightarrow\mathbb{R}$ that measures an extent to which datapoint $d\in\mathbf{R}^{c}$ contains a feature $g\in\mathbb{R}^{s}$.
+Suppose that we have already defined a function match: $(\mathbb{R}^{c},\mathbb{R}^{s})\rightarrow\mathbb{R}$ that measures an extent to which datapoint $d\in\mathbb{R}^{c}$ contains a feature $g\in\mathbb{R}^{s}$.
 Now let’s define
@@
 If $\mathbf{L}$ captures domain topology adequately then we may think of matching $f_{\mathbf{P}\mathbf{L}}$ as a form of local inhibition along $\mathbf{L}_{\mathbf{P}}(f)=\{L(p)(f),p\in P\}$ and in this sense semantic feature defines a XOR gate over $\mathbf{L}_{\mathbf{P}}(f)$.
-On the other hand every $f_{p}\in\mathbf{L}_{\mathbf{P}}(f)$ can be viewed as a conjunction (AND gate) of its non-zero coordinates666i.e.
+On the other hand, every $f_{p}\in\mathbf{L}_{\mathbf{P}}(f)$ can be viewed as a conjunction (AND gate) of its non-zero coordinates777i.e.
 $f_{p}$ is a conjunction of lower-level features..
-Thus semantic features turn out to be a natural way of expressing logical claims about the classical world - in the language that can be used by neural networks.
-This is particularly apparent in the context of logical semantic features which can express logical sentences explicitly, e.g.
+Thus, semantic features turn out to be a natural way of expressing logical claims about the classical world - in the language that can be used by neural networks.
+This is particularly apparent in the context of logical semantic features, which can express logical sentences explicitly, e.g.
 sentence $(A\vee B)\wedge\neg C\wedge\neg D$ can be expressed by a logical $f_{\mathbf{P}\mathbf{L}}$ where:
@@
   • for real-valued $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=\mathrm{int}(d==g)$ where $d,g\in\mathbb{R}$
-  • for convolutional777an alternative definition would treat every $g^{i}$ as a separate semantic feature; we stick to the max option for simplicity.
+  • for convolutional888an alternative definition would treat every $g^{i}$ as a separate semantic feature; we stick to the max option for simplicity.
 $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=\max_{0<=i<c_{out}}(d*\frac{g^{i}}{||g^{i}||_{2}})$ where $g$ is a 2D convolutional kernel of shape $(c_{out},c_{in},k,k)$, $g^{i}$ is the $i$-th convolutional filter of $g$, $d$ is an image patch of shape $(c_{in},k,k)$
-  • for affine $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=d\cdot\frac{g}{||g||_{2}}$ where $g$ is a 2D image of shape $(c_{in},k,k)$ and $d$ is a 2D image patch of the same shape
+  • for affine $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=d\cdot\frac{g}{||g||_{2}}$ where $g$ is a 2D image of shape $(c_{in},k,k)$ and $d$ is a 2D image patch of the same shape (treated as vectors in $\mathbb{R}^{c}$)
   • for logical $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=d\cdot g$ for $g,d\in\mathbb{R}^{c}$
@@
 Its basic building block is a pixel $x\in[0,1]$.
-Despite being allowed to assume any value between 00 and $1$ it is actually (semantically) a ternary object - it’s either ON, OFF or MEH888the MEH state is an undecided state in-between..
+Despite being allowed to assume any value between 00 and $1$ it is actually (semantically) a ternary object - it’s either ON, OFF or MEH999the MEH state is an undecided state in-between..
 Therefore the semantic space should squash the $[0,1]$ interval into those 3 values.
@@
 The layer performs $\mathrm{SFmatch}(d(x,y),f_{\mathbf{P}\mathbf{L}})$ with every pixel $(x,y)$ of image $d$.
-This means that it matches the two rotated filters $g^{0}$ and $g^{1}$ with the $5\times 5$ neighbourhood of $(x,y)$ and takes the maximum across all the matches999a single number, not a pair of numbers, as in the chosen definition of match for convolutional $f_{\mathbf{P}\mathbf{L}}$..
+This means that it matches the two rotated filters $g^{0}$ and $g^{1}$ with the $5\times 5$ neighbourhood of $(x,y)$ and takes the maximum across all the matches101010a single number, not a pair of numbers, as in the chosen definition of match for convolutional $f_{\mathbf{P}\mathbf{L}}$..
 Intuitively, this layer checks if pixels have adequately structured neighbourhoods.
@@
 In short they are shapes at locations.
-The semantic identity of shape is not affected by small101010close to the identity augmented matrix.
+The semantic identity of shape is not affected by small111111close to the identity augmented matrix.
 affine transformations.
@@
 The model was trained to optimize the CrossEntropyLoss for 15 epochs with Adam optimizer, learning rate $3e{-3}$, weight decay $3e{-6}$.
-The only augmentation was Gaussian noise $\sim\mathcal{N}(0.05,0.25)$ applied with $0.7$ probability111111mean is slightly positive because of the sparsity of data - subtracting pixel value is on average more harmful than adding it..
+The only augmentation was Gaussian noise $\sim\mathcal{N}(0.05,0.25)$ applied with $0.7$ probability121212mean is slightly positive because of the sparsity of data - subtracting pixel value is on average more harmful than adding it..
 Those parameters were chosen "by hand" after few rounds of experimentation.
@@
 Semantic features require locality engineering as they rely on the adequate domain-specific definition of the locality function.
-Therefore they won’t be applicable in contexts where such locality is hard to define.
-On the other hand it opens up interesting areas for research.
-The Logical Layer has been developed as an attempt to formalize the idea of how the generic hidden layer activation pattern should look like; however it might turn out to be prone to the "no free lunch" theorem (e.g.
+Therefore, they won’t be applicable in contexts where such locality is hard to define.
+On the other hand, it opens up interesting areas for research.
+The Logical Layer has been developed as an attempt to formalize the idea of how the generic hidden layer activation pattern should look like; however, it might turn out to be prone to the "no free lunch" theorem (e.g.
 on MNIST the properly-initialized linear probe seems to be sufficient).
@@
 8 Further research
-Future work would have to involve more complex datasets and more nuanced semantic features.
-For example notice the analogy between affine semantic features and the SIFT [21] algorithm - the learnable pose sampling may be interpreted as finding Keypoints and their Descriptors, while previous layers (Two Step and Convolutional Semantic) are essentially performing input normalization.
+Notice the analogy between affine semantic features and the SIFT [26] algorithm - the learnable pose sampling may be interpreted as finding Keypoints and their Descriptors, while the initial 2 layers of the PoC network essentially perform input normalization.
 This might be a helpful analogy for designing more advanced semantic features that sample across varying scales, sharpness or illumination levels.
 There might be a concern that adding more layers to the network would reduce its interpretability.
-However the role of additional layers would be to widen the receptive field of the network.
+However, the role of additional layers would be to widen the receptive field of the network.
 The Affine Layer defined earlier matches shapes-at-locations of center-aligned objects.
-Adding additional layer would involve a sliding window approach (or Keypoint Detection similar to that in SIFT) to match center-aligned objects inside the window and integrate this with the approximate global location of the window; then the Global Affine Layer matches objects-at-locations which form a scene (a higher-level object).
-This would in fact be similar to applying Affine Layer to the "zoomed-out" version of the input.
+Adding additional layer would involve a sliding window approach (or Keypoint Detection similar to that in SIFT) to match center-aligned objects inside the window and integrate this with the approximate global location of the window; then the "Global" Affine Layer would match objects-at-locations forming a scene (a higher-level object).
+This would in fact be similar to applying Affine Layer to the "zoomed-out" input.
 The ablations in Section 6.4 suggest that adversarial vulnerability of DNNs might be mostly due to the fragility of the initial layers.
-Therefore another direction forward would be to re-examine the known approaches to XAI and try to "bug fix" their encoder networks in similar fashion.
+Therefore, another direction forward would be to re-examine the known approaches to XAI and try to "bug fix" their encoder networks in similar fashion.
 Other ideas for further work on semantic features include:
   • use semantic features in self-supervised setting to obtain interpretable dimensionality reduction
-  • study non-affine perturbations of features, e.g. projective perturbations
+  • study non-affine perturbations of spatial features, e.g. projective perturbations
   • find adequate semantic topology for color space